OpenStudy (anonymous):

Solve x^4-34x^2=-225 (Factor the expression)

OpenStudy (anonymous):

You can let some variable a = x^2 Then rearrange your equation to get x^4 - 34x^2 + 225 = 0 From there, substitute in your value of a to get a^2 - 34a + 225 = 0 Now factorise like any normal quadratic. When you have your values of a, be sure to sub back in a = x^2 so that rather than having, for example, a = 9, you would have x^2 = 9 and therefore x = 3 or -3.